OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 7x - 2y = 21 4x + y = 57 21 -21 9 -9

OpenStudy (anonymous):

7x - 2y = 21 4x + y = 57 Multiply everything in the bottom equation by 2 and you get: 7x - 2y = 21 8x + 2y = 114 Then add both equations together and you get: 15x = 135 Then divide both sides by 15 and you get x = 9

OpenStudy (anonymous):

7x - 2y = 21---------- A 4x + y = 57 ---------- B -------------------------------- A + 2*B -------------------------------- 7x - 2y = 21 + 8x + 2y = 114 ------------- 15x =135 x = 9 --------------- put this value in eq B 4*9 + y = 57 y = 57-36 y=21